HUBBLE SPACE TELESCOPE DAILY REPORT #1790 PERIOD COVERED: 0000Z (UTC) 01/07/97 - 0000Z (UTC) 01/08/97 Daily Status Report as of 008/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FGS 6486 (Orbits of Pre-Main Sequence Binaries) The FGSs were used in the TRANS mode to make astrometric = observations of the per-main sequence, binary star systems ZZ-Tau and = DF-Tau in order to help compute its orbital parameters. The observations = will be used to make orbital maps of T-Tauri class stars. From these = orbital parameters they will be able to derive the first reliable mass = measurements for low mass pre-main sequence stars. The observations were = completed as scheduled, and no anomalies were reported. 1.2 Completed FGS 6922 (Blue TF Calibration) FGS#3 was used to make blue transfer function (TF) observations = through the pupil, clear, and neutral density filters at the center of = the field-of-view (FOV). These transfer function calibrations will be = used to deduce the presence of secondary stars in multiple systems and to = calibrate non-zero angular diamater of extended sources. The = observations were completed as planned, and no anomalies were noted. 1.3 Completed YSAFE (Safemode Recovery of the HST Science Intruments - = FOS) This step was executed as scheduled, and no anomalies were noted. 1.4 Completed KSAFE (Costar to Hold Mode) This step was executed as scheduled, and no anomalies were noted. 1.5 Completed XSAFE (Safemode Recovery of the HST Science Intruments - = FOC) This step was executed as scheduled, and no anomalies were noted. 1.6 Completed USAFE (WF/PC-2 Safe to Hold Mode) This step was executed as scheduled, and no anomalies were noted. 1.7 Completed FOS/BL 6781 (Soft X-Rays and the UV Spectra of a = Complete QSO Sample) The Faint Object Spectrograph (Blue Side) was used to obtain = ultraviolet spectra of the low-Z quasi-stellar object PG1427+480. The = ionization continuum of this object has been studied using the ROSAT = satellite. The strengths and line profiles of strong and weak = ultraviolet emission and absorption lines and the ultraviolet continuum = were studied. The observations were completed as planned and no anomalies= were reported. 1.8 Completed WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Operations Notes: Much of the operational day was spent returning to normal science = operations following the successful installation of the new NSSC-1 = software, a key step in preparation for the second servicing mission. 3.0 SIGNIFICANT FORTHCOMING EVENTS: SMGT-32 (SSR Re-Test) will be conducted today. Tomorrow, the 30-hour Joint Integrated Simulation #5 = (EVA4/Planning/Deploy) will begin at approximately 9 a.m. /DMH